OpenStudy (anonymous):

Which of the following graphs represents the function f(x) = -2x^3 - x^2 + 3x + 1

sammixboo (sammixboo):

Welcome to Open Study! Having any questions or problems? Just ask me or anyone with a purple A next to their username!

OpenStudy (ranga):

The end behavior of this function will be very much like that of -x^3. As x -> - inf, y-> +inf and as x -> + inf, y-> -inf

OpenStudy (anonymous):

I still don't understand lol

OpenStudy (ranga):

This is a cubic function. The last two graphs are more or less like a parabola and so you can rule them out. It is either the first or the second choice. Put x = 0, what is f(x)? How about x = -1 and x = +1. You can quickly find out which graph it is. f(x) = -2x^3 - x^2 + 3x + 1 f(0) = ? f(1) = ? f(-1) = ?

OpenStudy (anonymous):

Oh i understand what your saying now

OpenStudy (ranga):

cool. what is your answer?

OpenStudy (anonymous):

Wait so f(1) = -2x^3 - x^2 + 3x + 1 is -3 ? and f(-1) = = -2x^3 - x^2 + 3x + 1 is 7 ? so would the answer be B?

OpenStudy (ranga):

f(x) = -2x^3 - x^2 + 3x + 1 f(0) = 0 0 0 + 1 = 1 f(1) = -2(1)^3 - (1)^2 + 3(1) + 1 = -2 - 1 + 3 + 1 = 1 f(-1) = -2(-1)^3 - (-1)^2 + 3(-1) + 1 = +2 - 1 - 3 + 1 = -1

OpenStudy (anonymous):

OOOOOh wow i'm so bad at math im so sorry, it would be A then correct because it touches at (1,0) ?

OpenStudy (ranga):

We are looking for a curve that has the points: (0, 1), (1, 1) and (-1, -1)

OpenStudy (ranga):

X Y 0 1 1 1 -1 -1

OpenStudy (anonymous):

Oh okay I see what your saying, so it's B. so whenever I do problems like these I just look for the points on the graph the curves make? One more question, how do I know which number to plug into f(x) or is it just any numbers? Btw thanks a lot for helping me because math is my worst subject.

OpenStudy (ranga):

B is correct. x = 0, 1 and -1 are the EASIEST points to plug into a polynomial. In this problem, for example, f(x) = -2x^3 - x^2 + 3x + 1 It is so easy to find f(0). All the x terms will be 0 and you will be left with 1. So when x = 0, f(0) = 1 (that is, when x = 0, y = 1 on the graph). This one point alone is sufficient to rule out A, C and D and you can correctly conclude it is B.

OpenStudy (anonymous):

Thanks a bunch you really helped me a lot!

OpenStudy (ranga):

You are very welcome.